OpenStudy (anonymous):

HELP(: Find the least common multiple of x^3-x^2+x-1 and x^2-1 . Write the answer in factored form.

OpenStudy (anonymous):

First, factor the equation. You should get: (x^2 + 1) and (x+1) (x-1) Your end result should be: (x^2 +1)(x+1)(x-1)

jhonyy9 (jhonyy9):

x^3 -x^2 +x-1 =x^2(x-1) +x-1 =(x-1)(x^2 +1) x^2 -1 =(x-1)(x+1)